Consider frequency modulation with a sinusoidal message signal (tone modulation) m(t) = a cos(2πfmt). The modulated signal is: vFM(t) = Ac cos[2πfct + βf sin(2πfmt)]. with βf = kfa/fm being the modulation index. What is the bandwidth of the modulated signal.
Answer #1

S:

Given the modulated signal is

VFM(t) = Ac cos(2πfct + f sin(2πfmt))

Where f= kf.a/fm =  f/f​​​​​​​​​​m

  f - modulation index

fm - maximum frequency

  f - frequency deviation ( f= kf.a)

The bandwidth of the modulated signal is calculated by Carson's rule, which states that nearly all of the power of a frequency- modulated signal lies within the bandwidth B i,e

B = 2(f +f​​​​​​m) kHz (or) Hz

= 2fm(1+f)